# Efficient Neural Combinatorial Optimization Solver for the Min-max Heterogeneous Capacitated Vehicle Routing Problem

### v1

@@ IV ECHO
IV-A Node and Vehicle Encoders
@@ IV-A1 Node Encoder
where $j\in\{0,\cdots,N\}$, and the trainable parameters include $\bm{W}_{no}\in\mathbb{R}^{3\times d}$, $\bm{W}_{ed}\in\mathbb{R}^{N\times d}$ and the deport token $\text{DT}\in\mathbb{R}^{d}$. Following the prior study [21], we exploit a learnable depot token to distinguish the depot from customer nodes during training.
@@
where $\bm{W}_{1}\in\mathbb{R}^{d\times 4d}$, $\bm{W}_{2}\in\mathbb{R}^{4d\times d}$, $\bm{W}^{O}\in\mathbb{R}^{d\times d}$ and $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}\in\mathbb{R}^{d\times d^{\prime}}$ are trainable parameters, and the number of head is set to 8, i.e., $d^{\prime}=d/8$. The symbol ||||| | denotes the concatenation operator.
@@
where $\bm{W}_{g}\in\mathbb{R}^{2d\times 1}$. By adopting this design, ECHO efficiently integrates local topological features into the node embedding. It is worth mentioning that while certain studies [40, 41, 42] also highlight the importance of local topological relationships, ECHO departs fundamentally in its methodology. Specifically, our ECHO learns to aggregate the local topological features in the embedding space, whereas these prior studies reshape attention scores among nodes during decoding using these features.
IV-A2 Vehicle Encoder
Because each vehicle’s current location corresponds to a node’s coordinates, the vehicle embedding should inherently incorporate the associated node embedding. Therefore, the vehicle encoder must fuse these node embeddings with vehicle‑specific attributes to produce the vehicle embedding. To this end, we adopt the vehicle encoder architecture proposed by the prior study [21]. Specifically, similar to the node encoder, the vehicle encoder first embeds the vehicle attributes (i.e., $\chi_{i},\rho_{i},\hat{\rho}_{i}^{t},$ and $\delta^{t}_{i}$, see Table I for definitions about these attributes), using two fully connected layers with ReLU activations, as follows:
@@ IV-A2 Vehicle Encoder
IV-B Decoder of ECHO
To efficiently select vehicle-node pairs, we design a decoder inspired by the CLIP architecture [43]. While the prior study [21] also employed a similar CLIP-based architecture, our decoder incorporates fundamentally different structural components. In particular, we propose a Parameter-Free Cross Attention (PFCA) mechanism, which explicitly amplifies the influence of the node selected at the $t-1$th time step, thereby reducing myopic decision-making.
Specifically, given the vehicle embedding $\mathbf{M}^{t}\in\mathbb{R}^{M\times d}$ and node embedding $\mathbf{N}\in\mathbb{R}^{(N+1)\times d}$ produced by their respective encoders, the proposed decoder first integrates the vehicle embedding selected at the $t-1$th time step, $\mathbf{M}^{t-1}_{s}$, with the node embedding $\mathbf{N}$ to yield the updated node embedding $\hat{\mathbf{N}}^{t}$, computed as follows:
@@ IV-B Decoder of ECHO
Subsequently, the vehicle embedding $\mathbf{M}^{t}$ and the updated node embedding $\hat{\mathbf{N}}^{t}$ are combined to compute the attention scores for all vehicle-node pairs at the current time step. In addition, to ensure feasibility, we mask invalid pairs by setting them to $\infty$- ∞. For example, if a vehicle lacks sufficient remaining capacity to fulfill a node’s demand, its corresponding vehicle–node pair is masked. Formally, the attention score $\beta_{i,j}^{t}$ is defined as follows:
@@
where $\tanh(\cdot)$ denotes the hyperbolic tangent activation function, $\lambda$ denotes the hyper-parameter set identically to that in the prior study [21].
@@
After computing the probability distribution, ECHO constructs solutions for the given MMHCVRP instance using either Greedy or Sampling strategies (see Section IV-A for more details on these strategies).
IV-C Data Augment Method
@@
Data augment techniques have been widely applied to stabilize the RL-based training process of NCO solvers in canonical single-vehicle VRPs [44, 45]. However, in the single‑vehicle setting, these methods need only exploit the node symmetry of the VRP. In contrast, MMHCVRP exhibits both node symmetry and vehicle‐permutation invariance. To the best of our knowledge, no prior work has simultaneously considered both intrinsic properties in solving the MMHCVRP, which may underlie their suboptimal performance.
TABLE II: Node Augment Transformations
@@
[…]
As shown in Fig. 4, to address this open challenge, we propose a tailored data augment method that integrates vehicle permutation invariance and node symmetry. Specifically, following the node augment approach of prior study [44], we generate $K=8$ augmentations per instance by reflecting node coordinates across multiple axes, as shown in Table II. Concurrently, for each augmented instance, we sample a random vehicle permutation of the vehicle set. Note that, since the total number of vehicle permutations $M!$ may be smaller than $K$, duplicate permutations across augmentations are unavoidable (e.g., for $M=3,3!=6<8$).
We train ECHO using the well-known REINFORCE algorithm [46] with a shared baseline similar to the prior studies [44, 45]. As detailed in Section III-A, the objective of MMHCVRP is to minimize the maximum vehicle travel time. Accordingly, we define the total reward $R$ as the negative of this objective. The policy gradient is computed as follows:
@@
TABLE III: Performance Comparison of different Algorithms on HCVRP Instances across Varying Node Scales $N$ and Vehicle Numbers $M$
$N$ 60 100
$M$ 3 5 7 3 5 7
Metric Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%$\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$
SISR $\dagger$ 6.57 - 271 4.00 - 274 2.91 - 276 10.29 - 615 6.17 - 623 4.45 - 625
GA $\dagger$ 9.21 40.18 233 6.89 72.25 320 5.98 105.50 405 15.33 48.98 479 10.93 77.15 623 9.10 104.49 772
SA $\dagger$ 7.04 7.15 130 4.39 9.75 289 3.30 13.40 362 11.13 8.16 434 6.80 10.21 557 5.01 12.58 678
AM (ICLR’19) $\dagger$ (g.) 8.49 29.22 0.08 5.51 37.75 0.08 4.15 42.61 0.09 12.68 23.23 0.14 8.10 31.28 0.13 6.13 37.75 0.13
ET (AAAI’24) $\dagger$ (g.) 7.58 15.37 0.15 4.76 19.00 0.17 3.58 23.02 0.16 11.74 14.09 0.25 7.25 17.50 0.25 5.23 17.53 0.26
DPN (ICML’24) $\dagger$ (g.) 7.50 14.16 0.18 4.60 15.00 0.19 3.45 18.56 0.26 11.54 12.15 0.30 6.94 12.48 0.40 4.98 11.91 0.43
DRL (TCYB’22) $\dagger$ (g.) 7.43 13.09 0.19 4.71 17.75 0.22 3.60 23.71 0.25 11.44 11.18 0.32 7.06 14.42 0.37 5.38 20.90 0.43
2D-Ptr (AAMAS’24) (g.) 7.20 9.59 0.04 4.48 12.00 0.05 3.31 13.75 0.05 11.12 8.07 0.08 6.75 9.40 0.08 4.92 10.56 0.08
PARCO (arXiv’25) $\dagger$ (g.) 7.12 8.37 0.04 4.40 10.00 0.05 3.25 11.68 0.05 10.98 6.71 0.06 6.61 7.13 0.05 4.79 7.64 0.05
ECHO (ours) (g.) 6.83 3.93 0.05 4.18 4.60 0.06 3.05 4.71 0.06 10.68 3.82 0.07 6.40 3.74 0.09 4.61 3.61 0.09
AM (ICLR’19) $\dagger$ (s.) 7.62 15.98 0.14 4.82 20.50 0.13 3.63 24.74 0.14 11.82 14.87 0.29 7.45 20.75 0.28 5.58 25.39 0.28
ET (AAAI’24) $\dagger$ (s.) 7.14 8.68 0.21 4.46 11.50 0.22 3.33 14.43 0.22 11.20 8.84 0.41 6.85 11.02 0.38 4.98 11.91 0.40
DPN (ICML’24) $\dagger$ (s.) 7.08 7.76 0.25 4.35 8.75 0.28 3.20 9.97 0.38 11.04 7.29 0.48 6.66 7.94 0.52 4.79 7.64 0.78
DRL (TCYB’22) $\dagger$ (s.) 6.97 6.09 0.30 4.34 8.50 0.36 3.25 11.68 0.43 10.90 5.93 0.60 6.65 7.78 0.76 4.98 11.91 0.92
2D-Ptr (AAMAS’24) (s.) 6.82 3.81 0.07 4.20 5.00 0.08 3.09 6.19 0.09 10.71 4.08 0.13 6.46 4.70 0.16 4.68 5.17 0.18
PARCO (arXiv’25) $\dagger$ (s.) 6.82 3.81 0.05 4.17 4.25 0.05 3.06 5.15 0.07 10.61 3.11 0.08 6.36 3.08 0.08 4.58 2.92 0.09
ECHO (ours) (s.) 6.64 1.01 0.28 4.03 0.96 0.37 2.93 0.98 0.47 10.44 1.50 10.49 6.24 1.12 0.51 4.48 0.77 1.14
  • Note: Symbol $\dagger$ denotes that the results are taken from the prior study [23], while the rest are obtained in this work.
V Experimental Results
@@
Section V-A first describes datasets, baseline methods, and evaluation metrics. Section V-B then compares the performance of different NCO solvers. Next, Section V-C evaluates the generalization ability of ECHO across different scales and distribution patterns. Finally, Section V-D conducts ablation studies to assess the effectiveness of all proposed components.
V-A Experimental settings
V-A1 Datasets and Hardware configuration
@@
V-A2 Baselines and Evaluation Metrics
We compare ECHO against six NCO benchmarks, including AM [28], DRL [20], ET [39], DPN [22], 2D-Ptr [21], and PARCO [23]. Of these, DRL, 2D-Ptr, and PARCO are specifically designed for the MMHCVRP, ET and DPN target the Min-max TSP, and AM is tailored to TSP and CVRP. In addition, following the prior studies [21, 20], we also include the heuristic algorithms GA [47], SA [48], and the SOTA SISR [49] for comparison. The evaluation metrics include the objective value (Obj.), Gap, and computation time (Time), where Gap is defined as the relative difference in the ”objective value” between each method and the SISR heuristic. In addition, for NCO solvers, we report the results under two decoding strategies, namely Greedy and Sampling decoding strategies. With Greedy decoding, the solver always selects vehicle(s) and node(s) with the maximum probability at each time step. Whereas the sampling strategy randomly selects vehicle(s) and node(s) utilizing the learned probability, constructs the number of $k$ candidate solutions, and ultimately reports the best among them. In all experiments conducted in the paper, $k$ is set to 1280, following the precedent established by [28].
V-B Performance Comparison
In Table III, we compare the performance of ECHO against various heuristics and NCO solvers on MMHCVRP instances with varying numbers of vehicles ($M=3,5,7$) and nodes ($N=60,100$). As shown in Table III, ECHO achieves SOTA performance across varying numbers of vehicles and nodes compared with the other NCO solvers. When adopting the sampling decoding strategy, ECHO reduces the average gap to 3.77% and 2.67% relative to the AR-based 2D-Ptr and PAR-based PARCO solvers, respectively. The high-level performance reflects the combined contributions of all proposed components and methods (see Section V-D). Notably, even under greedy decoding, ECHO matches PARCO’s performance in gap achieved with sampling. Compared to the SOTA heuristic SISR, ECHO exhibits an average gap of approximately 1%, while running over 100× faster, making it more suitable for time-sensitive real-world scenarios.
V-C Generalization Studies
To assess the cross-scale and cross-distribution generalization abilities of ECHO, we conduct extensive experiments in this subsection. Specifically, in Figs. 5 and 6, we present the generalization performance of ECHO versus 2D-Ptr across varying node and vehicle scales, respectively. The abbreviation M3-N60 denotes a configuration with three vehicles and 60 nodes. In Fig. 7, we assess the cross-distribution generalization performance of ECHO and 2D-Ptr. To generate clustered test instances, following the procedure of the prior study [42], we sample three cluster centers for each instance, assign each node uniformly to one center, and perturb each node’s coordinates with Gaussian noise to induce the clustered distribution.
(a) 2D-Ptr
(b) ECHO

### v2

@@ 4 ECHO
4.1 Node and Vehicle Encoders
@@ 4.1.1 Node Encoder
where $j\in\{0,\cdots,N\}$, and the trainable parameters include $\bm{W}_{no}\in\mathbb{R}^{3\times d}$, $\bm{W}_{ed}\in\mathbb{R}^{N\times d}$ and the deport token $\text{DT}\in\mathbb{R}^{d}$. Following the prior study [2dptr], we exploit a learnable depot token to distinguish the depot from customer nodes during training.
@@
where $\bm{W}_{1}\in\mathbb{R}^{d\times 4d}$, $\bm{W}_{2}\in\mathbb{R}^{4d\times d}$, $\bm{W}^{O}\in\mathbb{R}^{d\times d}$ and $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}\in\mathbb{R}^{d\times d^{\prime}}$ are trainable parameters, and the number of head is set to 8, i.e., $d^{\prime}=d/8$. The symbol $||$ denotes the concatenation operator.
@@
where $\bm{W}_{g}\in\mathbb{R}^{2d\times 1}$. By adopting this design, ECHO efficiently integrates local topological features into the node embedding. It is worth mentioning that while certain studies [wang2024distance, reevo] also highlight the importance of local topological relationships, ECHO departs fundamentally in its methodology. Specifically, our ECHO learns to aggregate the local topological features in the embedding space, whereas these prior studies reshape attention scores among nodes during decoding using these features.
4.1.2 Vehicle Encoder
Because each vehicle’s current location corresponds to a node’s coordinates, the vehicle embedding should inherently incorporate the associated node embedding. Therefore, the vehicle encoder must fuse these node embeddings with vehicle‑specific attributes to produce the vehicle embedding. To this end, we adopt the vehicle encoder architecture proposed by the prior study [2dptr]. Specifically, similar to the node encoder, the vehicle encoder first embeds the vehicle attributes (i.e., $\chi_{i},\rho_{i},\hat{\rho}_{i}^{t},$ and $\delta^{t}_{i}$, see Table 1 for definitions about these attributes), using two fully connected layers with ReLU activations, as follows:
@@ 4.1.2 Vehicle Encoder
4.2 Decoder of ECHO
To efficiently select vehicle-node pairs, we design a decoder inspired by the CLIP architecture [clip]. While the prior study [2dptr] also employed a similar CLIP-based architecture, our decoder incorporates fundamentally different structural components. In particular, we propose a Parameter-Free Cross Attention (PFCA) mechanism, which explicitly amplifies the influence of the node selected at the t1t-1th time step, thereby reducing myopic decision-making.
Specifically, given the vehicle embedding $\mathbf{M}^{t}\in\mathbb{R}^{M\times d}$ and node embedding $\mathbf{N}\in\mathbb{R}^{(N+1)\times d}$ produced by their respective encoders, the proposed decoder first integrates the vehicle embedding selected at the t1t-1th time step, $\mathbf{M}^{t-1}_{s}$, with the node embedding $\mathbf{N}$ to yield the updated node embedding $\hat{\mathbf{N}}^{t}$, computed as follows:
@@ 4.2 Decoder of ECHO
Subsequently, the vehicle embedding $\mathbf{M}^{t}$ and the updated node embedding $\hat{\mathbf{N}}^{t}$ are combined to compute the attention scores for all vehicle-node pairs at the current time step. In addition, to ensure feasibility, we mask invalid pairs by setting them to -$\infty$. For example, if a vehicle lacks sufficient remaining capacity to fulfill a node’s demand, its corresponding vehicle-node pair is masked. Formally, the attention score $\beta_{i,j}^{t}$ is defined as follows:
@@
where $\tanh(\cdot)$ denotes the hyperbolic tangent activation function, $\lambda$ denotes the hyper-parameter set identically to that in the prior study [2dptr].
@@
After computing the probability distribution, ECHO constructs solutions for the given MMHCVRP instance using either Greedy or Sampling strategies (see Section 5.1 for more details on these strategies).
4.3 Data Augment Method
@@
Data augment techniques have been widely applied to stabilize the RL-based training process of NCO solvers in canonical single-vehicle VRPs [pomo, sym-nco]. However, in the single‑vehicle setting, these methods need only exploit the node symmetry of the VRP. In contrast, MMHCVRP exhibits both node symmetry and vehicle‐permutation invariance. To the best of our knowledge, no prior work has simultaneously considered both intrinsic properties in solving the MMHCVRP, which may underlie their suboptimal performance.
Table 2: Node Augment Transformations
@@
$(x,y)$ $(y,x)$
(x,1y)(x,1-y) (y,1x)(y,1-x)
(1x,y)(1-x,y) (1y,x)(1-y,x)
(1x,1y)(1-x,1-y) (1y,1x)(1-y,1-x)
As shown in Fig. 4, to address this open challenge, we propose a tailored data augment method that integrates vehicle permutation invariance and node symmetry. Specifically, following the node augment approach of prior study [pomo], we generate $K=8$ augmentations per instance by reflecting node coordinates across multiple axes, as shown in Table 2. Concurrently, for each augmented instance, we sample a random vehicle permutation of the vehicle set. Note that, since the total number of vehicle permutations $M!$ may be smaller than $K$, duplicate permutations across augmentations are unavoidable (e.g., for $M=3,3!=6<8$).
We train ECHO using the well-known REINFORCE algorithm [williams_simple_1992] with a shared baseline similar to the prior studies [pomo, sym-nco]. As detailed in Section 3.1, the objective of MMHCVRP is to minimize the maximum vehicle travel time. Accordingly, we define the total reward $R$ as the negative of this objective. The policy gradient is computed as follows:
@@
Table 3: Performance Comparison of different Algorithms on HCVRP Instances across Varying Node Scales $N=60,100$ with a Fixed Vehicle Number $M=3$
$M$ 3
NN 60 100
Metric Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$
SISR $\dagger$ 6.57 - 271 10.29 - 615
GA $\dagger$ 9.21 40.18 233 15.33 48.98 479
SA $\dagger$ 7.04 7.15 130 11.13 8.16 434
AM (ICLR’19) $\dagger$ (g.) 8.49 29.22 0.08 12.68 23.23 0.14
ET (AAAI’24) $\dagger$ (g.) 7.58 15.37 0.15 11.74 14.09 0.25
DPN (ICML’24) $\dagger$ (g.) 7.50 14.16 0.18 11.54 12.15 0.30
DRL (TCYB’22) $\dagger$ (g.) 7.43 13.09 0.19 11.44 11.18 0.32
2D-Ptr (AAMAS’24) (g.) 7.20 9.59 0.04 11.12 8.07 0.08
PARCO (NeurIPS’25) $\dagger$ (g.) 7.12 8.37 0.04 10.98 6.71 0.06
ECHO (ours) (g.) 6.83 3.93 0.05 10.68 3.82 0.07
AM (ICLR’19) $\dagger$ (s.) 7.62 15.98 0.14 11.82 14.87 0.29
ET (AAAI’24) $\dagger$ (s.) 7.14 8.68 0.21 11.20 8.84 0.41
DPN (ICML’24) $\dagger$ (s.) 7.08 7.76 0.25 11.04 7.29 0.48
DRL (TCYB’22) $\dagger$ (s.) 6.97 6.09 0.30 10.90 5.93 0.60
2D-Ptr (AAMAS’24) (s.) 6.82 3.81 0.07 10.71 4.08 0.13
PARCO (NeurIPS’25) $\dagger$ (s.) 6.82 3.81 0.05 10.61 3.11 0.08
ECHO (ours) (s.) 6.64 1.01 0.28 10.44 1.50 0.49
Table 4: Performance Comparison of different Algorithms on HCVRP Instances across Varying Node Scales $N=60,100$ with a Fixed Vehicle Number $M=5$
$M$ 5
NN 60 100
Metric Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$
SISR $\dagger$ 4.00 - 274 6.17 - 623
GA $\dagger$ 6.89 72.25 320 10.93 77.15 623
SA $\dagger$ 4.39 9.75 289 6.80 10.21 557
AM (ICLR’19) $\dagger$ (g.) 5.51 37.75 0.08 8.10 31.28 0.13
ET (AAAI’24) $\dagger$ (g.) 4.76 19.00 0.17 7.25 17.50 0.25
DPN (ICML’24) $\dagger$ (g.) 4.60 15.00 0.19 6.94 12.48 0.40
DRL (TCYB’22) $\dagger$ (g.) 4.71 17.75 0.22 7.06 14.42 0.37
2D-Ptr (AAMAS’24) (g.) 4.48 12.00 0.05 6.75 9.40 0.08
PARCO (NeurIPS’25) $\dagger$ (g.) 4.40 10.00 0.05 6.61 7.13 0.05
ECHO (ours) (g.) 4.18 4.60 0.06 6.40 3.74 0.09
AM (ICLR’19) $\dagger$ (s.) 4.82 20.50 0.13 7.45 20.75 0.28
ET (AAAI’24) $\dagger$ (s.) 4.46 11.50 0.22 6.85 11.02 0.38
DPN (ICML’24) $\dagger$ (s.) 4.35 8.75 0.28 6.66 7.94 0.52
DRL (TCYB’22) $\dagger$ (s.) 4.34 8.50 0.36 6.65 7.78 0.76
2D-Ptr (AAMAS’24) (s.) 4.20 5.00 0.08 6.46 4.70 0.16
PARCO (NeurIPS’25) $\dagger$ (s.) 4.17 4.25 0.05 6.36 3.08 0.08
ECHO (ours) (s.) 4.03 0.96 0.37 6.24 1.12 0.51
Table 5: Performance Comparison of different Algorithms on HCVRP Instances across Varying Node Scales $N=60,100$ with a Fixed Vehicle Number $M=7$
$M$ 7
NN 60 100
Metric Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$ Obj. $\downarrow$ Gap (%) $\downarrow$ Time (s) $\downarrow$
SISR $\dagger$ 2.91 - 276 4.45 - 625
GA $\dagger$ 5.98 105.50 405 9.10 104.49 772
SA $\dagger$ 3.30 13.40 362 5.01 12.58 678
AM (ICLR’19) $\dagger$ (g.) 4.15 42.61 0.09 6.13 37.75 0.13
ET (AAAI’24) $\dagger$ (g.) 3.58 23.02 0.16 5.23 17.53 0.26
DPN (ICML’24) $\dagger$ (g.) 3.45 18.56 0.26 4.98 11.91 0.43
DRL (TCYB’22) $\dagger$ (g.) 3.60 23.71 0.25 5.38 20.90 0.43
2D-Ptr (AAMAS’24) (g.) 3.31 13.75 0.05 4.92 10.56 0.08
PARCO (NeurIPS’25) $\dagger$ (g.) 3.25 11.68 0.05 4.79 7.64 0.05
ECHO (ours) (g.) 3.05 4.71 0.06 4.61 3.61 0.09
AM (ICLR’19) $\dagger$ (s.) 3.63 24.74 0.14 5.58 25.39 0.28
ET (AAAI’24) $\dagger$ (s.) 3.33 14.43 0.22 4.98 11.91 0.40
DPN (ICML’24) $\dagger$ (s.) 3.20 9.97 0.38 4.79 7.64 0.78
DRL (TCYB’22) $\dagger$ (s.) 3.25 11.68 0.43 4.98 11.91 0.92
2D-Ptr (AAMAS’24) (s.) 3.09 6.19 0.09 4.68 5.17 0.18
PARCO (NeurIPS’25) $\dagger$ (s.) 3.06 5.15 0.07 4.58 2.92 0.09
ECHO (ours) (s.) 2.93 0.98 0.47 4.48 0.77 1.14
5 Experimental Results
@@
To assess the performance of ECHO, we conduct extensive experiments in this section. Section 5.1 first describes datasets, baseline methods, and evaluation metrics. Section 5.2 then compares the performance of different NCO solvers. Next, Section 5.3 evaluates the generalization ability of ECHO across different scales and distribution patterns. Finally, Section 5.4 conducts ablation studies to assess the effectiveness of all proposed components.
5.1 Experimental Settings
Datasets and Hardware configuration: Following the prior studies [2dptr, drl], we assess the performance of ECHO on MMHCVRP instances with varying numbers of vehicles and nodes. Except for the cross-distribution generalization experiments in Section 5.3, the coordinates of customer nodes and depot node are uniformly sampled from the unit square $0,1]\times[0,1]$, and the demand of customer nodes are uniformly sampled from the set $\{1,2,\cdots,9\}$. For vehicles, capacities are uniformly sampled from $\{20,21,\cdots,40\}$, and speeds are uniformly sampled from the interval $[0.5,1]$. For each vehicle and node scale, the training dataset comprises the number of 1,280,000 instances, and the test dataset comprises 10,000 instances. All experiments are conducted on a computer equipped with an Intel(R) Core(TM) i9-12900K CPU and an NVIDIA RTX 4090 GPU (24GB).
Baselines and Evaluation Metrics: We compare ECHO against six NCO benchmarks, including AM [kool_attention_2019], DRL [drl], ET [et], DPN [dpn], 2D-Ptr [2dptr], and PARCO [parco]. Of these, DRL, 2D-Ptr, and PARCO are specifically designed for the MMHCVRP, ET and DPN target the Min-max TSP, and AM is tailored to TSP and CVRP. In addition, following the prior studies [2dptr, drl], we also include the heuristic algorithms GA [KARAKATIC2015519], SA [ILHAN2021100911], and the SOTA SISR [sisr] for comparison. The evaluation metrics include the objective value (Obj.), Gap, and computation time (Time), where Gap is defined as the relative difference in the ”objective value” between each method and the SISR heuristic. In addition, for NCO solvers, we report the results under two decoding strategies, namely Greedy and Sampling decoding strategies. With Greedy decoding, the solver always selects vehicle(s) and node(s) with the maximum probability at each time step. Whereas the sampling strategy randomly selects vehicle(s) and node(s) utilizing the learned probability, constructs the number of $k$ candidate solutions, and ultimately reports the best among them. In all experiments conducted in the paper, $k$ is set to 1280, following the precedent established by [kool_attention_2019].
5.2 Performance Comparison
@@
5.3 Generalization Studies
To assess the cross-scale and cross-distribution generalization abilities of ECHO, we conduct extensive experiments in this subsection. Specifically, in Figs. 5 and 6, we present the generalization performance of ECHO versus 2D-Ptr across varying node and vehicle scales, respectively. The abbreviation M3-N60 denotes a configuration with three vehicles and 60 nodes. In Fig. 7, we assess the cross-distribution generalization performance of ECHO and 2D-Ptr. To generate clustered test instances, following the procedure of the prior study [invit], we sample three cluster centers for each instance, assign each node uniformly to one center, and perturb each node’s coordinates with Gaussian noise to induce the clustered distribution.